# Understanding Long Videos with Multimodal Language Models

### v4

@@ Abstract
Large Language Models (LLMs) have allowed recent LLM-based approaches to achieve excellent performance on long-video understanding benchmarks. We investigate how extensive world knowledge and strong reasoning skills of underlying LLMs influence this strong performance. Surprisingly, we discover that LLM-based approaches can yield surprisingly good accuracy on long-video tasks with limited video information, sometimes even with no video specific information. Building on this, we explore injecting video-specific information into an LLM-based framework. We utilize off-the-shelf vision tools to extract three object-centric information modalities from videos, and then leverage natural language as a medium for fusing this information. Our resulting Multimodal Video Understanding (MVU) framework demonstrates state-of-the-art performance across multiple video understanding benchmarks. Strong performance also on robotics domain tasks establish its strong generality. Code: github.com/kahnchana/mvu
@@ 1 Introduction
In this work, we systematically study this question in the context of video question-answering (QnA) benchmarks, building two modality-constrained baselines to highlight our findings. These two frameworks are tagged Just-LLM and Single-Frame-VLM. The first is constrained to access only the task textual query (i.e. no task-specific visual information). The latter is given access to task context with an additional single center-frame from the video as input. We discover how these models perform significantly better than random prediction on multiple long-video understanding benchmarks (see Table 1, similar findings in Min et al. (2024)). In fact, the latter utilizing purely world knowledge and contextual information even outperforms multiple recent state-of-the-art video understanding works (see Table 2), challenging the notion of how much video information is actually utilized by existing approaches to solve these complex video QnA tasks.
@@
Motivated by our initial findings on modality-constrained performance, we study how to inject additional video-specific information into our framework using natural language in a concise and interpretable manner to further improve video understanding. We explore three forms of object-centric information modalities, develop pipelines requiring zero video-level training to extract such information using off-the-shelf vision tools, and utilize natural language to fuse this multi-modal information using templating operations. Our resulting approach, termed Multi-Modal Video Understanding (MVU) framework, while achieving state-of-the-art zero-shot performance across long-video understanding benchmarks, also exhibits better interpretability (e.g. exposing video-specific information utilized) through its language-based operation. Moreover, our proposed MVU exhibits generality with its strong performance even on robotics domain tasks.
@@
Figure 1: Overview of Framework: We propose three variants of our framework that solves complex long-video question-answering tasks. (left-top) Just-LLM utilizes only world knowledge with zero task-specific awareness. (left-bottom) Single-Frame-VLM processes an additional center frame to obtain task context but accesses no video specific information. (right) Our complete approach, MVU extracts three additional object-centric modalities followed by fusion in language space. LS refers to likelihood selection.
@@ 2 Related Work
Video Modality Exploration: Multiple recent works dissect video modality into individual components (Yun et al., 2022; Buch et al., 2022; Ranasinghe et al., 2021; Ramasinghe et al., 2018). Single frame baselines are one interesting sub-class (Buch et al., 2022; Davis & Bobick, 1997; Zhao et al., 2017; Safaei & Foroosh, 2019; Bilen et al., 2016). Extracting object-centric video modalities is another idea, spanning back to Davis & Bobick (1997) which extracts multiple small objects from frames followed by modeling relations across frames and objects. Similarly, Safaei & Foroosh (2019); Zhao et al. (2017) combine spatial information with single images to perform video tasks. However, these prior approaches focus on simple video tasks (i.e. action recognition) limited to visual modality. In contrast, our approach tackles the more complex language-tied task of long-video question answering that necessitates strong causal and temporal reasoning over long temporal windows. This task is also explored in Buch et al. (2022), but we differ with likelihood selection, multi-modal information fusion, and usage of modern LLMs.
@@
Long Video Question Answering: Long-video question-answering benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Early works explore querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019), followed by focus on temporal modeling of sequential events (Lei et al., 2018; 2020; Hosseini et al., 2022; Xiao et al., 2022a; b). While motivated by these works, MVU integrates such object information with large language models (LLMs) in a zero-shot manner requiring no video-level training. More recent works leverage LLMs (Yu et al., 2023b; Papalampidi et al., 2023; Wang et al., 2024b; Balavzevi’c et al., 2024; Wang et al., 2024a) to directly perform these tasks but require video-caption training. In contrast, our MVU operates zero-shot on these tasks requiring no video-level training. Zero-shot operation is explored in Wang et al. (2023a); Zhang et al. (2023a); Min et al. (2024); Wang et al. (2023b; 2024c), but we differ in using object-centric information modalities and efficient LLM sampling.
@@
Large Language Model Reasoning: Recent LLMs (OpenAI, 2023; Chowdhery et al., 2022; Chiang et al., 2023) demonstrate multiple forms of strong reasoning abilities (Kıcıman et al., 2023; Creswell & Shanahan, 2022; Liu et al., 2023b) including combining different information (Weston & Sukhbaatar, 2023). Their recent open-source variants (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2023) achieve equally promising skills using scaled-down models (Jiang et al., 2023) while also demonstrating strong world knowledge (Yu et al., 2023a; AlKhamissi et al., 2024; Zhao et al., 2023; Wang & Zhao, 2023; Xu et al., 2024; Li et al., 2023d) even in domains such as robotics (Li et al., 2024). In our work, we leverage these strengths of LLMs for complex video-language tasks, focused on disentangling the effect of their abilities for video QnA tasks.
@@
Language based Fusion: The idea of fusing different modality information using natural language as a medium has been explored in multiple recent works (Ranasinghe & Ryoo, 2023; Lin et al., 2023b; Hanu et al., 2022; Wang & Chen, 2017; Hanu et al., 2023; Zeng et al., 2022). In Ranasinghe & Ryoo (2023); Lin et al. (2023b), language is utilized as an implicit medium for self-supervising video action recognition. Multimodal information represented as language is fused with visual information for action recognition and robotics tasks in (Hanu et al., 2022; Wang & Chen, 2017; Hanu et al., 2023; Li et al., 2024). We utilize a similar language-as-a-medium fusion of multimodal information but explore this in the context of complex video-language tasks. Zeng et al. (2022) is most similar to our work, but we differ with focus on long-video tasks and object-centric information.
@@ 3.2 Likelihood Selection
Figure 2: Likelihood Selection Workflow: We illustrate how our proposed selection strategy can be efficiently parallelized and calculated with a simple cross-entropy loss in one forward pass, followed by an argmin operation, without iteratively generating multiple tokens.
@@ 5.1 Long Video Question Answering
5.2 Robotics Domain Action Recognition
We investigate generalization capabilities of our proposed MVU by evaluating across datasets from robotics domain Open X-Embodiment (Open-X-Embodiment-Collaboration et al., 2023), following a QnA style formulation of the dataset (details in Appendix D). We highlight visual differences of this data in Figure 4. We present evaluations in Table 4, which indicate clear performance improvements for MVU over the baseline from Zhang et al. (2023a). The purpose of this experiment is to evaluate the generality of our approach to video domains different from everyday natural videos. We take these promising results to indicate the strong generality of our framework. Furthermore, we note how our modality constrained variants do not perform significantly better than random on these robotics domain tasks (details in Appendix E). We attribute this to the significant domain shift in terms of the world of operation in this domain (i.e. robotics tasks tend to involve controlled environments very different to what humans face on an everyday basis).
@@ 5.3 Ablations
In this section, we systematically dissect our overall MVU framework to establish the usefulness of each of its individual component (see Appendix I for more ablations). We first ablate our three different information modalities and report these results in Table 5. Our evaluations indicate clear performance boosts from each of our object-centric information modalities.
@@
We next ablate our overall framework against the existing work, Zhang et al. (2023a), by replacing our MVU object-centric information pipeline with the frame description approach in Zhang et al. (2023a). We construct a setup identical to our framework except for the inputs to our final stage VLM replaced with frame level descriptions. These results reported in Table 7 indicate the clear significance and improvement of our proposed object-centric information pipeline over simple frame descriptions. The 8 frame variant is the same speed comparison as MVU uses captioner only on 8 frames. Our MVU outperforms both that and the slower 16 frame baseline. We also note the performance drop in the baseline when increasing the number of frames from 16 to 180. While consistent with observations in prior works for long-video tasks (Mangalam et al., 2023), we attribute this drop to decreased signal to noise ratio with introducing of additional frame descriptions. This further highlights the importance of selecting useful information from video frames and we reiterate how the object-centric information in our MVU framework serves this purpose.
@@ 6 Conclusion
We also presented two modality-constrained baselines that uncover surprising insights relevant to LLM based video QnA which serves as a basis for our subsequent MVU framework. Furthermore, these results highlight the need for careful evaluation of LLM-based video QnA approaches. Revisiting our original motivation on, “what we can learn from videos, beyond scene context understood from a single natural image”, in this work our two modality-constrained variants uncover surprising insights relevant to this question. We first achieve strong results on long-video understanding benchmarks using no video-specific data, and build over that baseline to showcase the additional performance gains achievable through injecting video-specific information.
@@
Our method utilizes multiple pretrained models, all of which are open-source with model weights freely available. We use the versions of these models hosted on HuggingFace https://huggingface.co for all our experiments. For LVNet baseline, we utilize code from https://github.com/jongwoopark7978/LVNet. We discuss all steps in our proposed algorithms in detail while also releasing relevant code. All evaluations we perform are on public datasets accessible by all (some behind evaluation servers to prevent test set contamination). Our code will also be released publicly.
@@ F.3 Detailed Prompting Example
Table A.6: Open-Ended Video QnA Evaluation: We present results on the ActivityNet dataset (Yu et al., 2019) that demonstrate strong performance of our proposed MVU framework. Accuracy (%) is reported. VT stands for video level training. We highlight how our MVU framework utilizes no video level training for any of its components and surpassed multiple approaches that rely on video-language training.
@@ Appendix G Open-Ended Video Question Answering
In this section, we explore the ability of our proposed MVU framework to operate on open-ended video question answering (QnA) tasks. For this purpose, we evaluate on the Activity-Net dataset (Yu et al., 2019) reporting the accuracy metric. We follow evaluation settings identical to Maaz et al. (2023) for these evaluations.

### v5

@@ Abstract
Large Language Models (LLMs) have allowed recent LLM-based approaches to achieve excellent performance on long-video understanding benchmarks. We investigate how extensive world knowledge and strong reasoning skills of underlying LLMs influence this strong performance. Surprisingly, we discover that LLM-based approaches can yield surprisingly good accuracy on long-video tasks with limited video information, sometimes even with no video-specific information. Building on this, we explore injecting video-specific information into an LLM-based framework. We utilize off-the-shelf vision tools to extract three object-centric information modalities from videos, and then leverage natural language as a medium for fusing this information. Our resulting Multimodal Video Understanding (MVU) framework demonstrates state-of-the-art performance across multiple video understanding benchmarks. Strong performance also on robotics domain tasks establishes its strong generality. Code: github.com/kahnchana/mvu
@@ 1 Introduction
In this work, we systematically study this question in the context of video question-answering (QnA) benchmarks, building two modality-constrained baselines to highlight our findings. These two frameworks are tagged Just-LLM and Single-Frame-VLM. The first is constrained to access only the task textual query (i.e. no task-specific visual information). The latter is given access to task context with an additional single center-frame from the video as input. We discover how these models perform significantly better than random prediction on multiple long-video understanding benchmarks (see Table 1, similar findings in Min et al. (2024)). In fact, the latter, utilizing purely world knowledge and contextual information, even outperforms multiple recent state-of-the-art video understanding works (see Table 2), challenging the notion of how much video information is actually utilized by existing approaches to solve these complex video QnA tasks.
@@
Motivated by our initial findings on modality-constrained performance, we study how to inject additional video-specific information into our framework using natural language in a concise and interpretable manner to further improve video understanding. We explore three forms of object-centric information modalities, develop pipelines requiring zero video-level training to extract such information using off-the-shelf vision tools, and utilize natural language to fuse this multi-modal information using templating operations. Our resulting approach, termed Multi-Modal Video Understanding (MVU) framework, while achieving state-of-the-art zero-shot performance across long-video understanding benchmarks, also exhibits better interpretability (e.g. exposing video-specific information utilized) through its language-based operation. Moreover, MVU exhibits generality with its strong performance even on robotics domain tasks (see zero-shot robot control in Appendix M).
@@
Figure 1: Overview of Framework: We propose three variants of our framework that solves complex long-video question-answering tasks. (left-top) Just-LLM utilizes only world knowledge with zero task-specific awareness. (left-bottom) Single-Frame-VLM processes an additional center frame to obtain task context but accesses no video specific information. (right) Our complete approach, MVU extracts three additional object-centric information modalities followed by fusion in language space. LS refers to likelihood selection.
@@ 2 Related Work
Video Modality Exploration: Multiple recent works dissect the video modality into individual components (Yun et al., 2022; Buch et al., 2022; Ranasinghe et al., 2021; Ramasinghe et al., 2018). Single frame baselines are one interesting sub-class (Buch et al., 2022; Davis & Bobick, 1997; Zhao et al., 2017; Safaei & Foroosh, 2019; Bilen et al., 2016). Extracting object-centric video modalities is another idea, spanning back to Davis & Bobick (1997) which extracts multiple small objects from frames followed by modeling relations across frames and objects. Similarly, Safaei & Foroosh (2019); Zhao et al. (2017) combine spatial information with single images to perform video tasks. However, these prior approaches focus on simple video tasks (i.e. action recognition) limited to visual modality. In contrast, our approach tackles the more complex language-tied task of long-video question answering that necessitates strong causal and temporal reasoning over long temporal windows. This task is also explored in Buch et al. (2022), but we differ with likelihood selection, multi-modal information fusion, and usage of modern LLMs.
@@
Long Video Question Answering: Long-video question-answering benchmarks are constructed to specifically test strong causal and temporal reasoning (Xiao et al., 2021) over long temporal windows (Mangalam et al., 2023). Early works explore querying objects or events based on referential and spatial relations (Xu et al., 2017; Zeng et al., 2017; Yu et al., 2019b), followed by focus on temporal modeling of sequential events (Lei et al., 2018; 2020; Hosseini et al., 2022; Xiao et al., 2022a; b). While motivated by these works, MVU integrates such object information with large language models (LLMs) in a zero-shot manner requiring no video-level training. More recent works leverage LLMs (Yu et al., 2023b; Papalampidi et al., 2023; Wang et al., 2024b; Balavzevi’c et al., 2024; Wang et al., 2024a) to directly perform these tasks but require video-caption training. In contrast, our MVU operates zero-shot on these tasks requiring no video-level training. Zero-shot operation is explored in Wang et al. (2023a); Zhang et al. (2023a); Min et al. (2024); Wang et al. (2023b; 2024c), but we differ in using object-centric information modalities and efficient LLM sampling.
@@
Large Language Model Reasoning: Recent LLMs (OpenAI, 2023; Chowdhery et al., 2022; Chiang et al., 2023) demonstrate multiple forms of strong reasoning abilities (Kıcıman et al., 2023; Creswell & Shanahan, 2022; Liu et al., 2023b) including combining different information (Weston & Sukhbaatar, 2023). Their recent open-source variants (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2023) achieve equally promising skills using scaled-down models (Jiang et al., 2023) while also demonstrating strong world knowledge (Yu et al., 2023a; AlKhamissi et al., 2024; Zhao et al., 2023; Wang & Zhao, 2023; Xu et al., 2024; Li et al., 2023d) even in domains such as robotics (Li et al., 2024b). In our work, we leverage these strengths of LLMs for complex video-language tasks, focused on disentangling the effect of their abilities for video QnA tasks.
@@
Language based Fusion: The idea of fusing different modality information using natural language as a medium has been explored in multiple recent works (Ranasinghe & Ryoo, 2023; Lin et al., 2023b; Hanu et al., 2022; Wang & Chen, 2017; Hanu et al., 2023; Zeng et al., 2022). In Ranasinghe & Ryoo (2023); Lin et al. (2023b), language is utilized as an implicit medium for self-supervising video action recognition. Multimodal information represented as language is fused with visual information for action recognition and robotics tasks in (Hanu et al., 2022; Wang & Chen, 2017; Hanu et al., 2023; Li et al., 2024b). We utilize a similar language-as-a-medium fusion of multimodal information, but explore this in the context of complex video-language tasks. Zeng et al. (2022) is most similar to our work, but we differ with focus on long-video tasks and object-centric information.
@@ 3.2 Likelihood Selection
Figure 2: Likelihood Selection Workflow: We illustrate how the likelihood selection strategy adapted for video QnA tasks can be efficiently parallelized (i.e. calculated with a simple cross-entropy loss in one forward pass, followed by an argmin operation), in contrast to the setting of iteratively generating multiple tokens.
@@ 5.1 Long Video Question Answering
5.2 Robotics Domain
Action Recognition: We investigate generalization capabilities of our proposed MVU by evaluating across datasets from robotics domain Open X-Embodiment, following a QnA style formulation of the dataset (details in Appendix D). We highlight visual differences of this data in Figure 4. We present evaluations in Table 4, which indicate clear performance improvements for MVU over the baseline from Zhang et al. (2023a). The purpose of this experiment is to evaluate the generality of our approach to video domains different from everyday natural videos. We take these promising results to indicate the strong generality of our framework. Furthermore, we note how our modality constrained variants do not perform significantly better than random on robotics domain tasks (details in Appendix E). We attribute this to significant domain shift in terms of the world of operation.
Robot Control: MVU extension for robot control in Appendix M further highlights our generality.
@@ 5.3 Ablations
In this section, we systematically dissect our overall MVU framework to establish the usefulness of each of its individual component (see Appendix I for more ablations). We first ablate our three different information modalities and report these results in Table 5. Our evaluations indicate clear performance improvements from each of our object-centric information modalities.
@@
We next ablate our overall framework against the existing work, Zhang et al. (2023a), by replacing our MVU object-centric information pipeline with the frame description approach in Zhang et al. (2023a). We construct a setup identical to our framework except for the inputs to our final stage VLM replaced with frame level descriptions. These results reported in Table 7 indicate the clear significance and improvement of our proposed object-centric information pipeline over simple frame descriptions. The 8 frame variant is the same speed comparison as MVU uses captioner only on 8 frames. Our MVU outperforms both that and the slower 16 frame baseline. We also note the performance drop in the baseline when increasing the number of frames from 16 to 180. While consistent with observations in prior works for long-video tasks (Mangalam et al., 2023), we attribute this drop to decreased signal-to-noise ratio with the introduction of additional frame descriptions. This further highlights the importance of selecting useful information from video frames, and we reiterate how the object-centric information in our MVU framework serves this purpose.
@@ 6 Conclusion
We also presented two modality-constrained baselines that uncover surprising insights relevant to LLM based video QnA which serves as a basis for our subsequent MVU framework. Furthermore, these results highlight the need for careful evaluation of LLM-based video QnA approaches. Revisiting our original motivation on “what we can learn from videos, beyond scene context understood from a single natural image”, in this work our two modality-constrained variants uncover surprising insights relevant to this question. We first achieve strong results on long-video understanding benchmarks using no video-specific data, and build over that baseline to showcase the additional performance gains achievable through injecting video-specific information.
@@
Our method utilizes multiple pretrained models, all of which are open-source with model weights freely available. We use the versions of these models hosted on HuggingFace https://huggingface.co for all our experiments. For state-of-the-art baseline LVNet, we utilize code from https://github.com/jongwoopark7978/LVNet. We discuss all steps in our proposed algorithms in detail while also releasing relevant code. All evaluations we perform are on public datasets accessible by all (some behind evaluation servers to prevent test set contamination). Our code is also publicly available at https://github.com/kahnchana/mvu.
@@ F.3 Detailed Prompting Example
Table A.6: Open-Ended Video QnA Evaluation: We present results on the ActivityNet dataset (Yu et al., 2019b) that demonstrate strong performance of our proposed MVU framework. Accuracy (%) is reported. VT stands for video level training. We highlight how our MVU framework utilizes no video level training for any of its components and surpassed multiple approaches that rely on video-language training.
@@ Appendix G Open-Ended Video Question Answering
In this section, we explore the ability of our proposed MVU framework to operate on open-ended video question answering (QnA) tasks. For this purpose, we evaluate on the Activity-Net dataset (Yu et al., 2019b) reporting the accuracy metric. We follow evaluation settings identical to Maaz et al. (2023) for these evaluations.
@@ Appendix I Additional Ablations
Appendix M Extension to Robot Control
We extend our MVU framework to robot manipulation tasks, as illustrated in Figure A.1. Inspired by prior work such as LLaRA (Li et al., 2024a), we formulate robot control—specifically, manipulation with a robot arm in tabletop environments—as a question-answering (QA) problem. This formulation allows us to directly leverage our MVU architecture, originally designed for video QA, to solve robotic manipulation tasks with minimal modification. We refer to this as MVU-R.
In contrast to the default MVU, our VLM here generates an estimate of the next gripper position, tagged Target Position Estimate. This is generated as an $(x,y)$ coordinate pair in image-space, motivated by prior work such as LLaRA (Li et al., 2024b). Paired with past object motion information and depth information, these coordinate sequences are processed by an LLM to provide a final gripper position. We provide a detailed example in Section M.1.
Figure A.1: MVU Robot Control Extension: We adapt MVU for robot manipulation (MVU-R) by framing control as a video question answering task, enabling zero-shot action prediction via vision-language prompting.
We conduct experiments in the MetaWorld (Yu et al., 2019a) simulated environment and evaluate on two manipulation tasks, following prior work such as AVDC (Ko et al., 2023b) and LTM (Ranasinghe et al., 2025). MetaWorld is a benchmark suite consisting of diverse robotic manipulation tasks performed by a simulated Sawyer robot arm in tabletop scenes. Each task provides access to egocentric RGB videos, object bounding boxes, and environment metadata.
To support 3D spatial understanding, we augment the input with depth maps from the simulator, similar to the AVDC baseline. In real-world applications, such depth information can be extracted using off-the-shelf depth estimation models (Ranasinghe et al., 2025; Ko et al., 2023b).
Table A.12 presents the results of our MVU-based robot control system. Remarkably, our model—without any robotics-specific fine-tuning—demonstrates promising zero-shot performance on these tasks, using only task-specific prompting. These findings suggest the potential of VLM/LLM-based models to serve as general-purpose, off-the-shelf agents for embodied control. We also emphasize how our MVU-R framework performs these tasks zero-shot while all prior works require some form of model training on expert task demonstrations (videos or action trajectories) from the environment.
Table A.12: MVU Robot Control Extension: We compare task success rates on two MetaWorld tasks: Door-Open and Door-Close. These tasks are selected following work in AVDC (Ko et al., 2023b) and LTM (Ranasinghe et al., 2025). Our MVU extension (MVU-R) achieves competitive zero-shot performance without any robotics-specific training, using only vision-language prompting. Results for baseline methods are taken from prior work, AVDC (Ko et al., 2023b) and LTM (Ranasinghe et al., 2025). While MVU-R achieves competitive results under zero-shot settings in simple object movement tasks (open / close door), it fails in complex object manipulation tasks. Tasks that require no gripper movement are marked “gripper ✗”.
Method Zero-Shot Task
Door-Open
(gripper ✗)
Door-Close
(gripper ✗)
Basket-Ball
(gripper ✓)
Shelf-Place
(gripper ✓)
BC-Scratch 21.3% 36.0% 0.0% 0.0%
BC-R3M 1.3% 58.7% 0.0% 0.0%
UniPi 0.0% 36.0% 0.0% 0.0%
AVDC 72.0% 89.3% 37.3% 18.7%
LTM 76.0% 94.7% 38.0% 15.2%
MVU (ours) 66.7% 77.3% 0.0% 0.0%
M.1 Implementation Details
Our MVU-R framework sequentially processes frames from a robot environment. At each time step, all objects are detected in current frame, connected to tracks in past frames, and processed by a VLM (both image and object motion tracks) to estimate a next gripper position. The VLM is prompted Based on provided motion information of all objects in scene, estimate next position of gripper. Only provide (x,y) coordinate.. This is used along with the MVU default system prompts that describe the motion trajectory format and specify additional video relevant information (even for robotics, we consider the past frame sequence as a video).
The next information fusion stage is similar to MVU, except for depth information and gripper position estimates. All object (x,y) coordinates are augmented with a depth $z$ coordinate obtained from a depth map of the image. This VLM output (after suitable templating operations to extract only x,y) is added as a gripper next position estimate. This information is processed by our LLM, similar to MVU, except in MVU-R a gripper position as $(x,y,z)$ is generated instead of an answer to a question.
M.2 Low Level Policy
As described previoisly, our MVU-R framework provides 3D coordinates, with $(x,y)$ based on image space and $z$ based on image-based depth map. We utilize the hand-crafted mapping of 3D locations in image-space to MetaWorld action vectors from AVDC (Ko et al., 2023b) and LTM (Ranasinghe et al., 2025) in our work. This also allows us fair and straightforward comparison to the evaluations from these prior works.
M.3 Limitations in MVU Robot Control
The key limitations of MVU-R are prompt sensitivity and lacking grasper awareness. When extending MVU-R to novel tasks, our system requires extensive hand-crafting of prompts. We consider this a trade-off for the zero-shot, training-free nature: our framework has never been trained on robotics tasks. MVU-R also currently cannot handle intelligent gripper manipulation (i.e. picking up objects). However, hand-crafting the gripper to hold when close to an object provides non-trivial performance on tasks, allowing MVU-R to perform selected grasping tasks. Similar findings are reported in AVDC (Ko et al., 2023b). In future work, we hope to extend gripper state (possibly extracted using segmentation modules) as an additional input to our VLM and LLM modules, providing MVU-R with more holistic scene awarenss.